# Formalization of Auslander–Buchsbaum–Serre criterion in Lean4

### v1

@@ 3.3. The Ischebeck Theorem
With the general definition moduleDepth, we can state the Ischebeck theorem as follows: for Noetherian local ring $R$ and non-trivial finitely generated module $M$ and $N$, $\operatorname{depth}(M,N)\geq\operatorname{depth}(M)-\dim(N)$W.
@@
[Module.Finite R M] [Nfin : Module.Finite R N] [Nontrivial M] [Nntr : Nontrivial N]
[Small.{v} R] : moduleDepth N M IsLocalRing.depth M -
@@
Since the case $\dim(N)=\infty$ is automatically correct, we prove the statement by strong induction on $\dim(N)$. If $\dim(N)=0$, then $\mathrm{Supp}(N)=V(\mathfrak{m})$, and hence $\operatorname{depth}(N,M)=\operatorname{depth}(M)$ by results established earlier in this section.
Now assume the statement holds for modules of dimension no greater than $n$, and let $\dim(N)=n+1$. In the development of associated primes one has an induction principle for finitely generated modules over a Noetherian ringW. It suffices to check the property in the following cases: $N=0$ (irrelevant here), $N\cong R/\mathfrak{p}$ for some prime ideal $\mathfrak{p}$, and that the property is preserved under short exact sequences. Preservation under exact sequences follows easily from Theorem 3.2 and a straightforward calculation. Be careful to apply the induction hypothesis appropriately to $N_{1}$ in the two subcases when $\dim(N_{1})\leq n$ and when $\dim(N_{1})=n+1$, because in this process the cited induction principle only provides hypotheses for modules of dimension $n+1$. The hypotheses in the former case come from the strong induction on dimension (and similarly for $N_{3}$). Thus the remaining and essential case is $N\cong R/\mathfrak{p}$, which is handled by the same argument as in [4, Lemma 15.2].
Using this theorem, we obtain a useful bound on $\operatorname{depth}$ by taking $N=R/\mathfrak{p}$ for $\mathfrak{p}\in\mathrm{Ass}(M)$. Since $\operatorname{depth}(M,R/\mathfrak{p})=0$, we have $\operatorname{depth}(M)\leq\dim(R/\mathfrak{p})$W. See also [4, Theorem 29].
@@
Moreover, from this we deduce the familiar inequalities $\operatorname{depth}(M)\leq\dim(M)\leq\dim(R)<\inftyW$.
@@ 4. Cohen–Macaulay Modules and Rings
Introduction:
In this section, we first give the definition of a Cohen–Macaulay module over a local ring. Then we prove that passing to the quotient by a regular element preserves the Cohen–Macaulay property, and that localization of a Cohen–Macaulay module is still Cohen–Macaulay. Similarly, we can establish the corresponding results for Cohen–Macaulay rings. In  4.1 we prove the unmixed theorem for Cohen–Macaulay rings; as a by‑product we obtain that a Cohen–Macaulay local ring is catenary. Since a polynomial ring over a Cohen–Macaulay ring is still Cohen–Macaulay (proved in  4.2), it will be straightforward to deduce that a Cohen–Macaulay ring is universally catenary once the preliminary constructions are in place.
With a well-developed definition of depth, we give the definition of a Cohen–Macaulay moduleW (over local ring) following [4, 16.A]. (In this section, let $R$ be a commutative ring, not assuming it to be Noetherian at first.)
@@ Introduction:
For a finitely generated Cohen–Macaulay module $M$ over a Noetherian local ring $R$, for any $\mathfrak{p}\in\mathrm{Ass}(M)$ we have $\mathrm{Ann}(M)\subseteq\mathfrak{p}$. As shown in  3.3, we know $\dim(M)=\operatorname{depth}(M)\leq\dim(R/\mathfrak{p})\leq\dim(R/\mathrm{Ann}(M))=\dim(M)$, so all inequalities are equalities. Hence $\dim(M)=\operatorname{depth}(M)=\dim(R/\mathfrak{p})$, forcing $\mathfrak{p}$ to be minimal prime over $\mathrm{Ann}(M)$. This yields the following two lemmasW W.
lemma depth_eq_dim_quotient_associated_prime_of_isCohenMacaulay (M : ModuleCat.{v} R)
[M.IsCohenMacaulay] [Module.Finite R M] [Nontrivial M] (p : Ideal R) (mem : p associatedPrimes R M) :
IsLocalRing.depth M = (ringKrullDim (R p)).unbot (quotient_prime_ringKrullDim_ne_bot mem.1)
lemma associated_prime_minimal_of_isCohenMacaulay (M : ModuleCat.{v} R)
[M.IsCohenMacaulay] [Module.Finite R M] [Nontrivial M] (mem : p associatedPrimes R M) :
p (Module.annihilator R M).minimalPrimes
Combining with the final result in  2.1, we know that the associated primes of a Cohen–Macaulay module are exactly the minimal primes over $\mathrm{Ann}(M)$W.
@@
From the results mentioned in  3.1 and  2.2, we know that for a finitely generated $R$-module $M$ with $R$ local, quotienting by an $M$-regular element reduces both the depth and the Krull dimension of $M$ by exactly one; therefore whether $M$ is Cohen–Macaulay is preserved under such a quotientW. A similar result holds for quotienting by a regular sequenceW. See also [4, Theorem 30(ii)]
lemma quotSMulTop_isCohenMacaulay_iff_isCohenMacaulay
(M : ModuleCat.{v} R) [Module.Finite R M] (r : R) (reg : IsSMulRegular M r) (mem : r maximalIdeal R) :
M.IsCohenMacaulay (ModuleCat.of R (QuotSMulTop r M)).IsCohenMacaulay
@@
We then formalized that localization of a Cohen–Macaulay module at prime ideal is still Cohen–MacaulayW.
@@
Here $R_{\mathfrak{p}}$ and $M_{\mathfrak{p}}$ denote localization of $R$ and $M$ at $\mathfrak{p}$ respectively.
To prove this proposition we first set up some lemmasW.
lemma isLocalization_at_prime_prime_depth_le_depth [IsLocalRing Rₚ] [Module.Finite R M]
[ntr : Nontrivial Mₚ] : p.depth M IsLocalRing.depth M
This uses the regular-sequence characterization of depth. As mentioned in  2.4, the image of an $M$-regular sequence contained in $\mathfrak{p}$ under the map $R\to R_{\mathfrak{p}}$ is $M_{\mathfrak{p}}$-regular, yielding the above inequality.
Since we know $\operatorname{depth}(M_{\mathfrak{p}})\leq\dim(M_{\mathfrak{p}})$ from  3.3, to prove $\operatorname{depth}_{\mathfrak{p}}(M)=\operatorname{depth}(M_{\mathfrak{p}})=\dim(M_{\mathfrak{p}})$, it remains to show $\operatorname{depth}_{\mathfrak{p}}(M)=\dim(M_{\mathfrak{p}})$W.
@@
From  3.3 we know that $\operatorname{depth}_{\mathfrak{p}}(M)\leq\operatorname{depth}(M_{\mathfrak{p}})$ is finite, so we proceed by induction on $\operatorname{depth}_{\mathfrak{p}}(M)$ following the proof in [4, Theorem 30(iii)].
Thus we finish proving that the localization at a prime of a Cohen–Macaulay module is Cohen–Macaulay, together with the equality $\operatorname{depth}_{\mathfrak{p}}(M)=\operatorname{depth}(M_{\mathfrak{p}})=\dim(M_{\mathfrak{p}})$.
The formalization of a Cohen–Macaulay local ring is done more directly rather than by extending the module notionW.
@@
This is equivalent to (ModuleCat.of R R).IsCohenMacaulay when $R$ is a local ringW.
From the result that localization of a Cohen–Macaulay module at a prime is Cohen–Macaulay, we obtain the corresponding version for Cohen–Macaulay local ringsW.
@@
In a straightforward way, we define a Cohen–Macaulay ring as a ring whose localization at every prime is a Cohen–Macaulay local ringW.
@@
Using the above localization result for Cohen–Macaulay local rings, we formalized that the Cohen–Macaulay property need only be verified at maximal idealsW, and that a Noetherian ring is Cohen–Macaulay if and only if its localizations at maximal ideals are Cohen–Macaulay local ringsW W.
@@
lemma IsCohenMacaulayRing.of_isCohenMacaulayLocalRing [IsCohenMacaulayLocalRing R] [IsNoetherianRing R] : IsCohenMacaulayRing R
lemma IsCohenMacaulayLocalRing.of_isLocalRing_of_isCohenMacaulayRing [IsLocalRing R] [IsNoetherianRing R] [IsCohenMacaulayRing R] : IsCohenMacaulayLocalRing R
We also establish the corresponding results about quotienting a ring by a regular element or a regular sequenceW W.
lemma quotient_span_regular_isCohenMacaulay_iff_isCohenMacaulay [IsLocalRing R] [IsNoetherianRing R]
(x : R) (reg : IsSMulRegular R x) (mem : x maximalIdeal R) :
IsCohenMacaulayLocalRing R IsCohenMacaulayLocalRing (R Ideal.span {x})
lemma quotient_regular_sequence_isCohenMacaulay_iff_isCohenMacaulay [IsLocalRing R] [IsNoetherianRing R]
(rs : List R) (reg : IsWeaklyRegular R rs) (mem : r rs, r maximalIdeal R) :
IsCohenMacaulayLocalRing R IsCohenMacaulayLocalRing (R Ideal.ofList rs)
4.1. Unmixed Theorem for Cohen–Macaulay Rings
@@
[4, Theorem 31(iii)]
We do not state the theorem as a single List.TFAE, instead we provide step-by-step lemmas, since some parts do not essentially use the Cohen–Macaulay or even the local hypothesisW W W. In the following lemmas we assume $R$ is Noetherian.
@@
The first two follow easily from Krull’s height theorem. The third follows from the fact that associated primes are minimal and that the quotient of a Cohen–Macaulay ring by a regular sequence is still Cohen–Macaulay, which we proved earlier in this section. The second and third together imply the following lemmaW.
@@
A by-product is the following lemmaW, i.e. [4, Theorem 31(i), Part1]:
@@
We then establish the following lemmaW to prove [4, Theorem 31(i), Part2]
lemma Ideal.primeHeight_add_ringKrullDim_quotient_eq_ringKrullDim [IsCohenMacaulayLocalRing R]
(p : Ideal R) [p.IsPrime] : p.primeHeight + ringKrullDim (R p) = ringKrullDim R
The formalization simply follows the proof in [4, Theorem 31(i), Part 2].
By simple manipulations of iSup and iInf we obtain the theoremW.
@@
From this result and the fact that localization of a Cohen–Macaulay ring is still Cohen–Macaulay, for prime ideals $\mathfrak{q}\subseteq\mathfrak{p}$ we have $\dim(R_{\mathfrak{p}})=0pt(\mathfrak{q}R_{\mathfrak{p}})+\dim(R_{\mathfrak{p}}/\mathfrak{q}R_{\mathfrak{p}})$, i.e. $0pt(\mathfrak{p})-0pt(\mathfrak{q})=0pt(\mathfrak{p}/\mathfrak{q})$. This implies that a Cohen–Macaulay local ring is catenary. However, since notions of catenary order and catenary rings are not yet in mathlib, we do not develop this further here.
To prove the unmixed theorem, we first define the notion of an unmixed ideal as followsW.
@@
With the results above, the formalization of the unmixed theorem follows easily by following the proof in [4, Theorem 32]W.
@@
4.2. Universal Catenary
Using the results proven in  4.1, to prove that Cohen–Macaulay rings are universally catenary it remains only to show that a polynomial ring over a Cohen–Macaulay ring is also Cohen–MacaulayW.

### v2

@@ 3.3. The Ischebeck Theorem
With the general definition moduleDepth, we can state the Ischebeck’s theorem (first introduced in [12]) as follows.W
Theorem 3.12 (Ischbeck).
For Noetherian local ring $R$ and non-trivial finitely generated module $M$ and $N$, $\operatorname{depth}(M,N)\geq\operatorname{depth}(M)-\dim(N)$
@@
[Module.Finite R M] [Nfin : Module.Finite R N] [Nontrivial M] [Nntr : Nontrivial N] [Small.{v} R] :
moduleDepth N M IsLocalRing.depth M -
@@
Proof.
Since the case $\dim(N)=\infty$ is not possible, we prove the statement by strong induction on $\dim(N)$. If $\dim(N)=0$, then $\mathrm{Supp}(N)=V(\mathfrak{m})$, and hence $\operatorname{depth}(N,M)=\operatorname{depth}(M)$ by Proposition  3.3.
Now assume the statement holds for modules of dimension no greater than $n$, and let $\dim(N)=n+1$. In the development of associated primes, one has an induction principle for finitely generated modules over a Noetherian ringW. It suffices to check the property in the following cases: $N=0$ (irrelevant here), $N\cong R/\mathfrak{p}$ for some prime ideal $\mathfrak{p}$, and that the property is preserved under extension by short exact sequences. Preservation under extension by short exact sequences follows easily from Lemma  3.5 and a straightforward calculation. Care must be taken to apply the induction hypothesis appropriately to $N_{1}$ in the two subcases when $\dim(N_{1})\leq n$ and when $\dim(N_{1})=n+1$, because in this process the cited induction principle only provides hypotheses for modules of dimension $n+1$. The hypotheses in the former case come from the strong induction on dimension (and similarly for $N_{3}$). Thus the remaining and essential case is $N\cong R/\mathfrak{p}$, which is handled by the same argument as in [16, Lemma 15.2]. ∎
Using this theorem, we obtain a useful bound on $\operatorname{depth}$.W
Corollary 3.13 ([16, Theorem 29]).
For finitely generated $R$-module $M$ and any $\mathfrak{p}\in\mathrm{Ass}(M)$, $\operatorname{depth}(M)\leq\dim(R/\mathfrak{p})$
@@
Proof.
Since $\operatorname{depth}(M,R/\mathfrak{p})=0$, by Theorem  3.12, we have $\operatorname{depth}(M)\leq\dim(R/\mathfrak{p})$. ∎
Moreover, from this we deduce the well-known inequalities $\operatorname{depth}(M)\leq\dim(M)\leq\dim(R)<\infty$.W
@@ 4. Cohen–Macaulay Modules and Rings
In this section, we first define Cohen–Macaulay modules over local rings. We then prove that passing to the quotient by a regular element preserves the Cohen–Macaulay property, and that localization of a Cohen–Macaulay module remains Cohen–Macaulay. Similarly, we establish corresponding results for Cohen–Macaulay rings. In Subsection  4.1, we prove the unmixedness theorem for Cohen–Macaulay rings; as a byproduct, we obtain that a Cohen–Macaulay local ring is catenary. Since a polynomial ring over a Cohen–Macaulay ring is again Cohen–Macaulay (proved in Subsection  4.2), it will be straightforward to deduce that a Cohen–Macaulay ring is universally catenary once the preliminary constructions are in place.
With a well-developed definition of depth, we define a Cohen–Macaulay module (over local ring) following [16, 16.A].W
Definition 4.1.
An $R$-module $M$ is a Cohen–Macaulay module if $M=0$ or $\operatorname{depth}(M)=\dim(M)$.
@@
Corollary  3.13 yields the following useful lemma, formalized as two separate statements.W W
Lemma 4.2 ([16, Theorem 30(i)]).
For $M$ a finitely generated Cohen–Macaulay module over Noetherian ring $R$, for any $\mathfrak{p}\in\mathrm{Ass}(M)$, $\dim(M)=\operatorname{depth}(M)=\dim(R/\mathfrak{p})$. Consequently $\mathfrak{p}$ is minimal prime over $\mathrm{Ann}(M)$.
Proof.
Since $\mathrm{Ann}(M)\subseteq\mathfrak{p}$, by Corollary  3.13, we have $\dim(M)=\operatorname{depth}(M)\leq\dim(R/\mathfrak{p})\leq\dim(R/\mathrm{Ann}(M))=\dim(M)$, so every inequality in the chain must be an equality. Hence $\dim(M)=\operatorname{depth}(M)=\dim(R/\mathfrak{p})$, forcing $\mathfrak{p}$ to be a minimal prime over $\mathrm{Ann}(M)$. ∎
Combining this with Lemma  2.2, we find that the associated primes of a Cohen–Macaulay module are exactly the minimal primes over $\mathrm{Ann}(M)$.W
@@
For a finitely generated $R$-module $M$ with $R$ local, quotienting by an $M$-regular element in the maximal ideal reduces both the depth and the Krull dimension of $M$ by exactly one; therefore, whether $M$ is Cohen–Macaulay is preserved under such a quotient.W
Lemma 4.3.
For $M$ a finitely generated module over $R$ and $x$ an $M$-regular element, then $M$ is Cohen–Macaulay if and only $M/xM$ is Cohen–Macaulay.
Proof.
This is a direct consequence of Theorem  3.6 and Lemma  2.3. ∎
A similar result holds for quotienting by a regular sequence, which would be used later in subsection  4.1.W
Lemma 4.4 ([16, Theorem 30(ii)]).
For $M$ a finitely generated module over $R$ and $(r_{1},\cdots,r_{k})$ an $M$-regular sequence, then $M$ is Cohen–Macaulay if and only $M/(r_{1},\cdots,r_{k})$ is Cohen–Macaulay.
@@
Proof.
This can be proved easily by induction on the length of the sequence. ∎
We then formalized that localization of a Cohen–Macaulay module at a prime ideal is again Cohen–Macaulay.
Lemma 4.5 ([16, Theorem 30(iii)]).
For $M$ a finitely generated Cohen–Macaulay module over $R$ and $\mathfrak{p}$ a prime ideal of $R$, then $M_{\mathfrak{p}}$ is Cohen–Macaulay over $R_{\mathfrak{p}}$. Furthermore, $\operatorname{depth}_{\mathfrak{p}}(M)=\operatorname{depth}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}})$.
This result is formalized as the following two separate statements.WW
@@
Proof.
To prove this result, we first set up a lemma.W
Proposition 4.6.
For $M$ a finitely generated module over $R$ and $\mathfrak{p}$ a prime ideal of $R$, $\operatorname{depth}_{\mathfrak{p}}(M)\leq\operatorname{depth}_{R_{\mathfrak{p}}}(M_{\mathfrak{p}})$.
Proof.
Here we use the regular sequence characterization of depth. As mentioned in Subsection  2.4, the image of an $M$-regular sequence contained in $\mathfrak{p}$ under the map $R\to R_{\mathfrak{p}}$ is $M_{\mathfrak{p}}$-regular, yielding the above inequality. ∎
Returning to the main proof, since we know $\operatorname{depth}(M_{\mathfrak{p}})\leq\dim(M_{\mathfrak{p}})$ from Subsection  3.3, to prove $\operatorname{depth}_{\mathfrak{p}}(M)=\operatorname{depth}(M_{\mathfrak{p}})=\dim(M_{\mathfrak{p}})$, it remains to show $\operatorname{depth}_{\mathfrak{p}}(M)=\dim(M_{\mathfrak{p}})$.
From subsection  3.3, we know that $\operatorname{depth}_{\mathfrak{p}}(M)\leq\operatorname{depth}(M_{\mathfrak{p}})$ is finite, so we proceed by induction on $\operatorname{depth}_{\mathfrak{p}}(M)$ following the proof in [16, Theorem 30(iii)]. This completes the proof that the localization at a prime of a Cohen–Macaulay module is Cohen–Macaulay, together with the equality $\operatorname{depth}_{\mathfrak{p}}(M)=\operatorname{depth}(M_{\mathfrak{p}})=\dim(M_{\mathfrak{p}})$. ∎
The formalization of a Cohen–Macaulay local ring is done more directly rather than by extending the module notion.W
Definition 4.7.
A local ring $R$ is a Cohen–Macaulay local ring if $\operatorname{depth}(R)=\dim(R)$.
@@
This is equivalent to (ModuleCat.of R R).IsCohenMacaulay when $R$ is a local ring.W
From the result that localization of a Cohen–Macaulay module at a prime is Cohen–Macaulay, we obtain the corresponding version for Cohen–Macaulay local rings as a corollary.W
Corollary 4.8.
For a Cohen–Macaulay local ring $R$ and $\mathfrak{p}$ a prime ideal of $R$, then $R_{\mathfrak{p}}$ is Cohen–Macaulay.
@@
We also establish the corresponding results of Lemma  4.3 and Lemma  4.4 for Cohen–Macaulay local rings.WW
Corollary 4.9.
For a local ring $R$ and an $R$-regular element $x$, $R$ is Cohen–Macaulay if and only if $R/(x)$ is Cohen–Macaulay.
Corollary 4.10.
For a local ring $R$ and an $R$-regular sequence $(r_{1},\cdots,r_{k})$, $R$ is Cohen–Macaulay if and only if $R/(r_{1},\cdots,r_{k})$ is Cohen–Macaulay.
With Definition  4.7, we can define general (non-local) Cohen–Macaulay rings.W
Definition 4.11.
A ring $R$ is a Cohen–Macaulay ring if $R_{\mathfrak{p}}$ is a Cohen–Macaulay local ring for every prime ideal $\mathfrak{p}$.
@@
Using the above localization result for Cohen–Macaulay local rings, we find that the Cohen–Macaulay property need only be verified at maximal ideals.W
Lemma 4.12.
A Noetherian ring $R$ is Cohen–Macaulay if and only if $R_{\mathfrak{m}}$ is Cohen–Macaulay local ring for every maximal ideal $\mathfrak{m}$.
@@
As a corollary, a Cohen–Macaulay local ring is a Cohen–Macaulay ringW and a local ring that is a Cohen–Macaulay ring is a Cohen–Macaulay local ring W.
4.1. Unmixedness Theorem for Cohen–Macaulay Rings
The unmixedness theorem for Cohen–Macaulay rings was first stated in the graded case by Macaulay in [13]. Here we follow the more modern formulation originating from [17].
@@ 4.1. Unmixedness Theorem for Cohen–Macaulay Rings
We do not state the theorem as a single List.TFAE, but instead provide step-by-step lemmas, since some parts do not essentially use the Cohen–Macaulay or even the local hypothesis.W W W In the following lemmas, we assume $R$ is Noetherian.
Proposition 4.14 (Theorem  4.13 $(1)\rightarrow(2)$).
For ring $R$ and a weakly regular sequence $(r_{1},\cdots,r_{k})$ generating a proper ideal, then $0pt(r_{1},\cdots,r_{k})=k$ .
@@
Proof.
The result follows easily by induction on the length of the sequence using Krull’s height theorem and the fact that a regular element is not contained in any minimal prime. ∎
Proposition 4.15 (Theorem  4.13 $(3)\rightarrow(4)$).
For local ring $R$ and a sequence $(r_{1},\cdots,r_{k})$, if $0pt(r_{1},\cdots,r_{k})=k$ then it can be extended to a system of parameters.
@@
Proof.
This result follows by adding elements to the tail inductively while ensuring that the height of the ideal generated by the sequence remains equal to the length of the sequence. ∎
Proposition 4.16 (Preparation for Theorem  4.13 $(4)\rightarrow(1)$).
For a Cohen–Macaulay local ring $R$, a system of parameters $(r_{1},\cdots,r_{k})$ form a regular sequence.
@@
Proof.
If the first $k$ elements $(r_{1},\cdots,r_{k})$ form a regular sequence, then by Corollary  4.10, $R/(r_{1},\cdots,r_{k})$ is Cohen–Macaulay. As all associated primes of $R/(r_{1},\cdots,r_{k})$ are minimal by Lemma  4.2, the fact that $r_{k+1}$ is not contained in any minimal prime of $R/(r_{1},\cdots,r_{k})$ implies it is a regular element. ∎
Propositions  4.15 and  4.16 together imply the following result.W
Proposition 4.17 (Theorem  4.13 $(4)\rightarrow(1)$).
For local ring $R$ and a sequence $(r_{1},\cdots,r_{k})$, if $0pt(r_{1},\cdots,r_{k})=k$ then it is a regular sequence.
@@
A by-product of Theorem  4.13 is the following lemma.W
Lemma 4.18 ([16, Theorem 31(i), Part1]).
For $R$ a Cohen–Macaulay local ring and $I$ a proper ideal of it, then $\operatorname{depth}_{I}(R)=0pt(I)$.
@@
Proof.
By Theorem  4.13, any regular sequence in $I$ has length at most $0pt(I)$, so the left-hand side is not greater than the right-hand side. Then as we can find $J\subseteq I$ such that $0pt(J)=0pt(I)$ and $J$ can be generated by $0pt(I)$ elements, Theorem  4.13 shows the generators form a regular sequence, then the right-hand side is not greater than the left-hand side. ∎
We then establish the following theorem which implies that a Cohen–Macaulay local ring is catenary. W
Theorem 4.19 ([16, Theorem 31(i), Part2]).
For $R$ a Cohen–Macaulay local ring and $I$ a proper ideal, we have $0pt(I)+\dim(R/I)=\dim(R)$.
@@
Proof.
Since minimal primes over $I$ is finite, by simple manipulations of iSup and iInf we can reduce to the case where $I$ is a prime ideal.W The formalization then follows the proof in [16, Theorem 31(i), Part 2]. ∎
From this result and the fact that localization of a Cohen–Macaulay ring is still Cohen–Macaulay, for prime ideals $\mathfrak{q}\subseteq\mathfrak{p}$, we have $\dim(R_{\mathfrak{p}})=0pt(\mathfrak{q}R_{\mathfrak{p}})+\dim(R_{\mathfrak{p}}/\mathfrak{q}R_{\mathfrak{p}})$, i.e. $0pt(\mathfrak{p})-0pt(\mathfrak{q})=0pt(\mathfrak{p}/\mathfrak{q})$. This implies that a Cohen–Macaulay local ring is catenary. However, since notions of catenary order and catenary rings are not yet in mathlib, we do not develop this further here.
Returning to the main theorem, we first define the notion of an unmixed ideal.W
Definition 4.20.
For $I$ an ideal of a Noetherian ring $R$, II is unmixed if for any $\mathfrak{p}\in\mathrm{Ass}(R/I)$, $0pt(\mathfrak{p})=0pt(I)$.
@@
In particular, if $I=(r_{1},\cdots,r_{k})$ with $0pt(I)=k$, then $I$ is unmixed if and only if $R/I$ has no embedded primes.
The unmixedness theorem can then be stated as the follows.
Theorem 4.21.
A Noetherian ring $R$ is Cohen–Macaulay if and only if for any ideal $I=(r_{1},\cdots,r_{k})$ with $0pt(I)=k$, II is unmixed.
@@
Proof.
With the results earlier in this section, the formalization of the unmixedness theorem follows easily from the proof in [16, Theorem 32].W
4.2. Universally Catenary
Using the results proven in Subsection  4.1, to prove that Cohen–Macaulay rings are universally catenary, it remains only to show that a polynomial ring over a Cohen–Macaulay ring is also Cohen–Macaulay.W